# Collusion-Resilience in Transaction Fee Mechanism Design

### v2

@@ 8.1 Static Revelation Principle: Bidding Rules That Output Single Bid
For proving $\Pi^{\prime}$ satisfies MIC, suppose $\Pi^{\prime}$ is not MIC for the sake of contradiction. In this case, there exists a bid vector ${\mathbf{b}}$ and a block $\tilde{\mathbf{B}}$ such that if the miner creates the block $\tilde{\mathbf{B}}$ instead of $\mathbf{I}^{\prime}({\mathbf{b}})$, the miner’s utility increases. Then, consider another miner under the mechanism $\Pi$, and imagine that the bid vector is $\beta({\mathbf{b}})$. Notice that $\mu(\beta(\tilde{\mathbf{B}}))=\mu^{\prime}(\tilde{\mathbf{B}})$ and $\mu(\mathbf{I}(\beta({\mathbf{b}})))=\mu^{\prime}(\mathbf{I}^{\prime}({\mathbf{b}}))$. Thus, if the miner creates the block $\beta(\tilde{\mathbf{B}})$ instead of $\mathbf{I}(\beta({\mathbf{b}}))$, the miner’s utility increases. It violates the fact that $\Pi$ is MIC. Thus, $\Pi^{\prime}$ must satisfy MIC.
@@ 8.2 Static Revelation Principle: Allowing Bidding Rules that Output Multiple Bids
To handle this difficulty, we need to relax the syntax of the inclusion rule. Earlier in our definitions (Section 2.1), we require that the inclusion rule outputs a subset of the input bid vector. In this section, we slightly relax this syntax requirement, and allow the honest inclusion rule to append some metadata for each bid. [^11] We assume the metadata do not occupy the block space. See Remark 3 for more discussion. The purpose of the metadata in the honest execution is to encode the auxiliary information for the blockchain’s confirmation/payment rules. This relaxation in the syntax is non-essential for the impossibility results in this paper, since all the proofs for our impossibility results (Sections 4, 5 and 6.2) still hold even if the honest inclusion rule can append the metadata. Now, we are ready to present the revelation principle.
@@ Appendix A Comparison of Collusion-Resilience Notions
{mdframed}
Even Auction
Parameters: the block size is infinite.
Mechanism:
  • Bidding rule. Each user submits two bids $(v,0)$, where $v$ is the true value.
  • Inclusion rule. Include all pending bids.
  • Confirmation rule. If the number of the included bids is odd, no one is confirmed. Otherwise, if the number of the included bids is even, all included bids are confirmed.
  • Payment rule and miner revenue rule. Everyone pays nothing, and the miner revenue is zero.
@@
Notice that whether the individually rational bidding strategy $\sigma$ outputs a single bid or multiple bids is important. If we require $\sigma$ to only output a single real as we defined in Definition 6, there is no non-trivial TFM that can satisfy UIC, MIC, and OCA-proofness as we proved in Theorem 6.9. However, if we allow $\sigma$ to output multiple bids, we obtain a feasibility in Section 7.2.
When we analyze whether a TFM satisfies $c$-SCP for some $c$, $1$-SCP is the weakest requirement since $c^{\prime}$-SCP always implies $c$-SCP for any $c^{\prime}>c$ by definition. Interestingly, even so, $1$-SCP is incomparable with global SCP and OCA-proofness. The relations between $1$-SCP, global SCP, and OCA-proofness is depicted in Figure 1.
@@
  • global SCP $\overset{\triangle}{\implies}$ OCA-proofness: As we discussed above, if the honest bidding rule only outputs a single real, global SCP implies OCA-proofness; if the honest bidding rule outputs multiple reals, Lemma A.1 states that global SCP does not imply OCA-proofness.
@@
{mdframed}
Pay-Nothing Auction
Parameters: the block size $k$.
Mechanism:
  • Bidding rule. Each user submits the true value.
  • Inclusion rule. Choose the highest up to $k$ bids.
  • Confirmation rule, payment rule and miner revenue rule. All included bids are confirmed, everyone pays nothing, and the miner revenue is zero.
@@
{mdframed}
Discount Auction
Parameters: the fixed price $r$, and the block size is infinite.
Mechanism:
  • Bidding rule. Each user submits the true value.
  • Inclusion rule. Include all pending bids.
  • Confirmation rule. Let $S$ be the set of all included bids, and define $f$ as
    $f(t)=\left\{\begin{matrix}r,&\text{if }t\leq 10;\\ r/2,&\text{otherwise}.\end{matrix}\right.$
    Let $t^{*}$ be the largest integer such that there are $t^{*}$ bids in $S$ larger than $f(t^{*})$. All bids larger than $f(t^{*})$ are confirmed.
  • Payment rule. Each confirmed bid pays $f(t^{*})$. Unconfirmed bids pay nothing.
  • Miner revenue rule. The miner revenue is zero, and all payments are burnt.
@@
Now, we show that the mechanism does not satisfy OCA-proofness. Imagine a scenario where there are $10$ users each with the true value $r$. Without injecting any fake bid, the lowest possible payment for each user is $f(10)=r$, so the social welfare is at most 00. However, if the global coalition inject one fake bid $r$, everyone’s payment will further lower to $f(11)=r/2$. In this case, the social welfare becomes $010r-11r/2=9r/2>$ 0. Thus, the mechanism does not satisfy OCA-proofness.

### v3

@@ 9.1 Static Revelation Principle: Bidding Rules That Output Single Bid
[…]
@@ 9.2 Static Revelation Principle: Allowing Bidding Rules that Output Multiple Bids
To handle this difficulty, we need to relax the syntax of the inclusion rule. Earlier in our definitions (Section 2.1), we require that the inclusion rule outputs a subset of the input bid vector. In this section, we slightly relax this syntax requirement, and allow the honest inclusion rule to append some metadata for each bid. [^13] We assume the metadata do not occupy the block space. See Remark 9.5 for more discussion. The purpose of the metadata in the honest execution is to encode the auxiliary information for the blockchain’s confirmation/payment rules. This relaxation in the syntax is non-essential for the impossibility results in this paper, since all the proofs for our impossibility results (Sections 4, 5 and 6.2) still hold even if the honest inclusion rule can append the metadata. Now, we are ready to present the revelation principle.
@@ Appendix A Comparison of Collusion-Resilience Notions
Even Auction Parameters: the block size is infinite. Mechanism: Bidding rule. Each user submits two bids $(v,0)$, where $v$ is the true value. Inclusion rule. Include all pending bids. Confirmation rule. If the number of the included bids is odd, no one is confirmed. Otherwise, if the number of the included bids is even, all included bids are confirmed. Payment rule and miner revenue rule. Everyone pays nothing, and the miner revenue is zero.
@@
Notice that whether the individually rational bidding strategy $\sigma$ outputs a single bid or multiple bids is important. If we require $\sigma$ to only output a single real as we defined in Definition 2.6, there is no non-trivial TFM that can satisfy UIC, MIC, and OCA-proofness as we proved in Theorem 6.10. However, if we allow $\sigma$ to output multiple bids, we obtain a feasibility in Section 7.2.
When we analyze whether a TFM satisfies $c$-SCP for some $c$, 11-SCP is the weakest requirement since $c^{\prime}$-SCP always implies $c$-SCP for any $c^{\prime}>c$ by definition. Interestingly, even so, $1$-SCP is incomparable with global SCP and OCA-proofness. The relations between $1$-SCP, global SCP, and OCA-proofness is depicted in Figure 1.
@@
[…]
@@
Pay-Nothing Auction Parameters: the block size $k$. Mechanism: Bidding rule. Each user submits the true value. Inclusion rule. Choose the highest up to $k$ bids. Confirmation rule, payment rule and miner revenue rule. All included bids are confirmed, everyone pays nothing, and the miner revenue is zero.
@@
Discount Auction Parameters: the fixed price $r$, and the block size is infinite. Mechanism: Bidding rule. Each user submits the true value. Inclusion rule. Include all pending bids. Confirmation rule. Let $S$ be the set of all included bids, and define $f$ as $f(t)=\left\{\begin{matrix}r,&\text{if }t\leq 10;\\ r/2,&\text{otherwise}.\end{matrix}\right$. Let $t^{*}$ be the largest integer such that there are $t^{*}$ bids in $S$ larger than $f(t^{*})$. All bids larger than $f(t^{*})$ are confirmed. Payment rule. Each confirmed bid pays $f(t^{*})$. Unconfirmed bids pay nothing. Miner revenue rule. The miner revenue is zero, and all payments are burnt.
@@
Now, we show that the mechanism does not satisfy OCA-proofness. Imagine a scenario where there are $10$ users each with the true value $r$. Without injecting any fake bid, the lowest possible payment for each user is $f(10)=r$, so the social welfare is at most 0. However, if the global coalition inject one fake bid $r$, everyone’s payment will further lower to $f(11)=r/2$. In this case, the social welfare becomes 10r11r/2=9r/2>010r-11r/2=9r/2>0. Thus, the mechanism does not satisfy OCA-proofness.
@@
Appendix B Alternative Proof of Theorem 5.6
The proof in Section 5 relies on the fact that the miner can deviate from the inclusion rule and choose an arbitrary subset of transactions to include in the block. In particular, Lemma 5.2 assumes that the global coalition can internally simulate the inclusion rule and strategically choose which transactions are included. In this section, we present an alternative proof of Theorem 5.6 which does not rely on deviations from the inclusion rule. As a result, this alternative proof rules out the possibility of achieving UIC, MIC, and global SCP even when the inclusion rule is strictly enforced, for example using trusted hardware.
The two proofs, however, are incomparable. The alternative proof in this section relies on the value domain being unbounded, since it analyzes the behavior of payments and burning as user values tend to infinity (see the proof intuition below). In contrast, the proof in Section 5 still holds even when the value domain is bounded, say $[0,M]$, assuming the confirmation probability under a single bid $a<M$ is non-zero. This assumption is mild. If it does not hold, i.e., if a single bid $a$ has positive confirmation probability only when $a=M$, then its payment must be $M$ by Myerson’s lemma. Consequently, the single user’s utility is always zero. Moreover, when the value domain is unbounded, we can always find a sufficiently large value $M$ such that confirmation probability is positive as shown in Section 5. Because each proof has its advantages, we include both in this paper.
Proof intuition.
Consider first a lonely world with only a single user submitting a true value $v$. By Myerson’s lemma, the user’s payment is given by $p(v)=v\cdot x(v)-\int_{0}^{v}x(t)\,dt$. We show that $\lim_{v\rightarrow\infty}p(v)/v=0$ (Lemma B.4), i.e., when the bid $v$ is sufficiently large, the user’s payment becomes negligible compared to $v$. Since the miner’s revenue is upper bounded by the total payment, the miner’s revenue is also negligible in the lonely world.
Now, consider a crowded world where $n$ users each submit the same bid $v$. We show that for any $n$, if the mechanism satisfies global SCP, then $\lim_{v\rightarrow\infty}Q_{n}(v)/v=0$, where $Q_{n}(v)$ denotes the total amount of burnt coin (Lemma B.3 and Lemma B.4). Conceptually, burnt coins represent the “payment” of the global coalition, so the proof parallels the analysis of a single user’s payment.
Next, we show that for any $n$, the miner revenue $\mu_{n}(v)$ in the crowded world must satisfy $\lim_{v\rightarrow\infty}\mu_{n}(v)/v=0$ (Lemma B.5). If this is not true, i.e., $\mu_{n}(v)$ is large with respect to $v$, then there must be a user $i$ whose payment is at least $\mu_{n}(v)/n$, which is also large with respect to $v$. However, as we have seen, the miner revenue is negligible in the lonely world. Thus, if there is only one user $i$ submitting $v$, the miner can inject n1n-1 fake bids, inflating user $i$’s payment to $\mu_{n}(v)/n$. Because the amount of burnt coins is negligible, the miner can extract $\mu_{n}(v)/n$ for user $i$, which is much larger than the honest case.
Finally, by global SCP, we show that adding an additional user to the system does not affect the joint utility of the original players (Lemma B.7). As we have shown, the miner’s revenue is negligible, so Lemma B.7 implies that each original user’s utility almost remains unchanged when an additional user is added. Thus, starting from the lonely world, where the only user has positive utility $u_{v}$, the total social welfare in the crowded world becomes $nu_{v}$, growing linearly with $n$. However, because the block size is finite, this is impossible.
Remark B.1 (On the permissionless setting).
In this alternative proof, it is essential to consider the permissionless setting, where the number of users is not fixed a priori and strategic players may inject an arbitrary number of fake bids. This assumption is critical because the proof does not rely on any deviation from the inclusion rule. Hence, it remains valid even if the miner is forced to follow the inclusion rule exactly.
If, in contrast, we further assume that the number of users is fixed a priori, the miner cannot inject any fake bids, and its strategy space becomes empty. In such a restricted setting, MIC would hold trivially. However, the second-price auction satisfies both UIC and global SCP, which implies that the impossibility does not hold.
The rest of this section is devoted to the formal proof. For convenience, we will use the following notation. Consider any number $n$ of users each with the true value $v$, and let $\mathbf{v}_{n}=(\underbrace{v,\dots,v}_{n})$. Let $X_{n}(v)=\sum_{i}x_{i}(\mathbf{v}_{n})$ denote the sum of all users’ allocation probability, $P_{n}(v)=\sum_{i}p_{i}(\mathbf{v}_{n})$ denote the total payment, $\mu_{n}(v)=\mu(\mathbf{v}_{n})$ denote the miner revenue, and $Q_{n}(v)=P_{n}(v)-\mu_{n}(v)$ denote the total burnt coin.
Lemma B.2 (Technical lemma implied by the proof of Myerson’s Lemma [Mye81, Har]).
Let $f(z)$ be a non-decreasing function. Suppose that $y\cdot(f(z)-f(y))\leq g(z)-g(y)\leq z\cdot(f(z)-f(y))$ for any $z\geq y\geq 0$, and moreover, $g(0)=0$. Then, it must be that
$g(z)=z\cdot f(z)-\int_{0}^{z}f(t)dt.$
Lemma B.3.
For any $n$, if the mechanism satisfies global SCP, then $X_{n}(\cdot)$ must be non-decreasing, and the function $Q_{n}(\cdot)$ satisfies
$Q_{n}(v)=v\cdot X_{n}(v)-\int_{0}^{v}X_{n}(t)dt.$
Proof.
If all users bid $b$ instead of their true value $v$, the social welfare is $v\cdot X_{n}(b)-Q_{n}(b)$. Consider any $0\leq y\leq z$, by global SCP, it must be $y\cdot X_{n}(y)-Q_{n}(y)\geq y\cdot X_{n}(z)-Q_{n}(z)$, and $z\cdot X_{n}(z)-Q_{n}(z)\geq z\cdot X_{n}(y)-Q_{n}(y)$. Thus, we obtain the following “burning sandwich” inequality:
$y\cdot[X_{n}(z)-X_{n}(y)]\leq Q_{n}(z)-Q_{n}(y)\leq z\cdot[X_{n}(z)-X_{n}(y)],$
which implies that $X_{n}(\cdot)$ must be non-decreasing. Because $p_{i}(0)=0$ for any user $i$, we have $Q_{n}(0)=0$. By Lemma B.2, we have
$Q_{n}(v)=v\cdot X_{n}(v)-\int_{0}^{v}X_{n}(t)dt.$
Lemma B.4.
Consider any natural number $k$ and any non-decreasing function $f(z):\mathbb{R}_{\geq 0}\rightarrow[0,k]$. We define function $g(\cdot)$ as
$g(v)=v\cdot f(v)-\int_{0}^{v}f(t)dt.$
Then, it holds that $\lim_{v\rightarrow\infty}g(v)/v=0$.
Proof.
It is equivalent to show that
$\lim_{v\rightarrow\infty}\left[f(v)-\frac{1}{v}\int_{0}^{v}f(t)dt\right]=0.$
It is well known that a bounded monotonic function must have a finite limit at infinity. Thus, we have $\lim_{v\rightarrow\infty}f(v)=L$ for some $L\in[0,k]$.
Next, we will show that $\lim_{v\rightarrow\infty}\frac{1}{v}\int_{0}^{v}f(t)dt=L$. In other words, we need to show that for any $\delta>0$, we can find a real number $v^{*}$ such that $\frac{1}{v}\int_{0}^{v}f(t)dt\in[L-\delta,L+\delta]$ for all $v>v^{*}$. Because $\lim_{v\rightarrow\infty}f(v)=L$, there exists a real number $v_{0}$ such that $f(v)\in[L-\delta,L+\delta]$ for all $v>v_{0}$, so we have
$(L-\delta)\cdot(v-v_{0})\leq\int_{v_{0}}^{v}f(t)dt\leq(L+\delta)\cdot(v-v_{0}).$
Because $\lim_{v\rightarrow\infty}(v-v_{0})/v=1$, we have
$\lim_{v\rightarrow\infty}\frac{1}{v}\int_{v_{0}}^{v}f(t)dt\in[L-\delta,L+\delta].$
We can split the integral as
$\frac{1}{v}\int_{0}^{v}f(t)dt=\frac{1}{v}\int_{0}^{v_{0}}f(t)dt+\frac{1}{v}\int_{v_{0}}^{v}f(t)dt.$
The term $\int_{0}^{v_{0}}f(t)dt$ is a constant independent of $v$, so we have $\lim_{v\rightarrow\infty}\frac{1}{v}\int_{0}^{v_{0}}f(t)dt=0$. Combining the argument above, we have
$L-\delta\leq\lim_{v\rightarrow\infty}\frac{1}{v}\int_{0}^{v}f(t)dt=\lim_{v\rightarrow\infty}\frac{1}{v}\int_{0}^{v_{0}}f(t)dt+\lim_{v\rightarrow\infty}\frac{1}{v}\int_{v_{0}}^{v}f(t)dt\leq L+\delta.$
Because the argument above holds for any $\delta>0$, we have $\lim_{v\rightarrow\infty}\frac{1}{v}\int_{0}^{v}f(t)dt=L$. Therefore, we have
$\lim_{v\rightarrow\infty}\left[f(v)-\frac{1}{v}\int_{0}^{v}f(t)dt\right]=0,$
so we conclude the proof. ∎
Lemma B.5.
Suppose the mechanism satisfies UIC, MIC, and global SCP. Then, for any $n$, it must be $\lim_{v\rightarrow\infty}\mu_{n}(v)/v=0$.
Proof.
Consider any case where there are $n$ users all submitting $v$. Because the total payment is $P_{n}(v)$, there must exist a user whose payment is at least $P_{n}(v)/n$.
Now, imagine that there is only one user with identity $i$ submitting $v$, so the miner’s utility is $\mu_{1}(v)$ in the honest execution. The miner randomly chooses n1n-1 fake identities $\mathcal{I}$, and checks whether $i$’s expected payment $p_{i}(\mathbf{v}_{n})$ is at least $P_{n}(v)/n$ given $n$ identities $\mathcal{I}\cup\{i\}$ all submitting $v$. If not, the miner plays honestly without injecting any fake bids. Otherwise, the miner injects n1n-1 fake bids $v$ on behalf of $\mathcal{I}$. In this case, $i$’s payment becomes $p_{i}(\mathbf{v}_{n})$, and the miner’s utility is $p_{i}(\mathbf{v}_{n})-Q_{n}(v)$. By MIC, it must be
$\mu_{1}(v)\geq p_{i}(\mathbf{v}_{n})-Q_{n}(v)\geq\frac{P_{n}(v)}{n}-Q_{n}(v).$ (3)
Because the honest miner revenue is upper bounded by the payment, we have $\mu_{1}(v)\leq P_{1}(v)$ and $\mu_{n}(v)\leq P_{n}(v)$. Together with Equation 3, we have
$\mu_{n}(v)\leq n\cdot[P_{1}(v)+Q_{n}(v)].$ (4)
Next, by UIC and Myerson’s lemma, the only user $i$’s payment is subject to $P_{1}(v)=v\cdot x(v)-\int_{0}^{v}x(v)dt$. Moreover, by Lemma B.3, we have $Q_{n}(v)=v\cdot X_{n}(v)-\int_{0}^{v}X_{n}(t)dt$. Thus, by Lemma B.4, we have $\lim_{v\rightarrow\infty}P_{1}(v)/v=0$ and $\lim_{v\rightarrow\infty}Q_{n}(v)/v=0$. Because the miner revenue must be non-negative, fix any $n$, Equation 4 implies $\lim_{v\rightarrow\infty}\mu_{n}(v)/v=0$. ∎
Next, we will show that for any user $i$, no matter how user $i$ bids, the joint utility of the rest of the world does not change. Formally, given a bid vector $\mathbf{b}$} and a user $i$, let $\mathsf{SW}({\mathbf{b}})$ be social welfare, i.e., the sum of the utilities of all users and the miner. Let $\mathsf{SW}_{-i}({\mathbf{b}})=\mathsf{SW}({\mathbf{b}})-\mathsf{util}_{i}({\mathbf{b}})$, i.e., the joint utility of the miner and all users except user $i$.
Lemma B.6.
Consider any (possibly randomized) TFM satisfying UIC and global SCP. Then, for any bid vector ${\mathbf{b}}_{-i}$ of an arbitrary length, for any user $i$, and for any $b_{i},b^{\prime}_{i}$, it holds that
$\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i})=\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b^{\prime}_{i}).$
Proof.
Because the mechanism satisfies UIC, user $i$’s payment is subject to Myerson’s lemma. Thus, given any bid vector ${\mathbf{b}}_{-i}$ from other users, user’s payment is $p_{i}({\mathbf{b}}_{-i},b_{i})=b_{i}\cdot x_{i}({\mathbf{b}}_{-i},b_{i})-\int_{0}^{b_{i}}x_{i}({\mathbf{b}}_{-i},t)dt$. Suppose user $i$’s true value is $b_{i}$. For any $b^{\prime}_{i}$, if user $i$ bids $b^{\prime}_{i}$ instead of $b_{i}$, user $i$’s loss can be represented as
$\mathsf{util}_{i}({\mathbf{b}}_{-i},b_{i})-\mathsf{util}_{i}({\mathbf{b}}_{-i},b^{\prime}_{i})$ $=[b_{i}\cdot x_{i}({\mathbf{b}}_{-i},b_{i})-p_{i}({\mathbf{b}}_{-i},b_{i})]-[b_{i}\cdot x_{i}({\mathbf{b}}_{-i},b^{\prime}_{i})-p_{i}({\mathbf{b}}_{-i},b^{\prime}_{i})]$
$\leq[b^{\prime}_{i}-b_{i}][x_{i}({\mathbf{b}}_{-i},b^{\prime}_{i})-x_{i}({\mathbf{b}}_{-i},b_{i})].$
Next, By global SCP, for any bid vector ${\mathbf{b}}_{-i}$, it must be
$\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i})+\mathsf{util}_{i}({\mathbf{b}}_{-i},b_{i})\geq\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b^{\prime}_{i})+\mathsf{util}_{i}({\mathbf{b}}_{-i},b^{\prime}_{i}).$
Notice that the argument above holds for any $b_{i}$ and $b^{\prime}_{i}$. Thus, for any $b_{i}$ and $b^{\prime}_{i}$, we have
$\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b^{\prime}_{i})-\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i})\leq[b^{\prime}_{i}-b_{i}][x_{i}({\mathbf{b}}_{-i},b^{\prime}_{i})-x_{i}({\mathbf{b}}_{-i},b_{i})].$ (5)
Now, we can divide the interval $b_{i},b^{\prime}_{i}]$ into $L$ equally sized segments $b_{i}^{(0)},\dots,b_{i}^{(L)}$ where $b_{i}^{(0)}=b_{i}$ and $b_{i}^{(L)}=b^{\prime}_{i}$. By Equation 5, we have
$\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b^{\prime}_{i})-\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i})$ $=\sum_{j=0}^{L-1}\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i}^{(j+1)})-\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i}^{(j)})$
$\leq\sum_{j=0}^{L-1}(b_{j}^{(j+1)}-b_{j}^{(j)})\cdot\left[x_{i}({\mathbf{b}}_{-i},b_{i}^{(j+1)})-x_{i}({\mathbf{b}}_{-i},b_{i}^{(j)})\right]$
$=\frac{b^{\prime}_{i}-b_{i}}{L}\cdot\left[x_{i}({\mathbf{b}}_{-i},b^{\prime}_{i})-x_{i}({\mathbf{b}}_{-i},b_{i})\right].$
By taking the limit for $L\rightarrow\infty$, we have $\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b^{\prime}_{i})-\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i})\leq 0$. Because the argument above holds for any $b_{i}$ and $b^{\prime}_{i}$, it must be $\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i})=\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b^{\prime}_{i})$. ∎
Lemma B.7.
Consider any (possibly randomized) TFM satisfying UIC and global SCP. Then, for any bid vector ${\mathbf{b}}_{-i}$ of an arbitrary length, for any user $i$, and for any $b_{i}$, it holds that
$\mathsf{SW}_{-i}({\mathbf{b}}_{-i})=\mathsf{SW}_{-i}({\mathbf{b}}_{-i},b_{i}).$
Proof.
First, we show that $\mathsf{SW}_{-i}({\mathbf{b}}_{-i})=\mathsf{SW}_{-i}({\mathbf{b}}_{-i},0_{i})$. To see this, if $\mathsf{SW}_{-i}({\mathbf{b}}_{-i})<\mathsf{SW}_{-i}({\mathbf{b}}_{-i},0_{i})$, then consider a world where users’ true values are ${\mathbf{b}}_{-i}$. The global coalition can inject a fake bid $0_{i}$, and the social welfare increases. On the other hand, if $\mathsf{SW}_{-i}({\mathbf{b}}_{-i})>\mathsf{SW}_{-i}({\mathbf{b}}_{-i},0_{i})$, then consider a world where users’ true values are $({\mathbf{b}}_{-i},0_{i})$. Because user $i$’s utility is zero under an honest execution, we have $\mathsf{SW}_{-i}({\mathbf{b}}_{-i},0_{i})=\mathsf{SW}({\mathbf{b}}_{-i},0_{i})$. If user $i$ drops out, the social welfare becomes $\mathsf{SW}_{-i}({\mathbf{b}}_{-i})$, which is larger than the honest execution. Thus, it must be $\mathsf{SW}_{-i}({\mathbf{b}}_{-i})=\mathsf{SW}_{-i}({\mathbf{b}}_{-i},0_{i})$. The lemma directly follows from the argument above and Lemma B.6. ∎
Theorem B.8.
No non-trivial, possibly randomized truthful TFM can simultaneously satisfy UIC, MIC, and global SCP (all in the ex post sense) when the block size is finite.
Proof.
For the sake of contradiction, suppose the mechanism is non-trivial; that is, there exists a scenario ${\mathbf{b}}=(b_{1},\dots,b_{t})$ such that some user $i$ has positive allocation probability $p_{i}({\mathbf{b}})$. We first show that for a sufficiently large $v_{0}$, the allocation probability under a single bid $v_{0}$ is positive. Let $v_{0}$ be any real number such that $p_{i}({\mathbf{b}})\cdot v_{0}>\sum_{j=1}^{t}b_{j}$. Then, imagine that there is only one user with true value $v_{0}$. If the allocation probability $x(v_{0})$ is zero, the global coalition can replace the user’s primary bid with $b_{i}$, inject the fake bids ${\mathbf{b}}_{-i}$, and pretend that the bid vector is $\mathbf{b}$}. This way, the user would get positive allocation probability $p_{i}({\mathbf{b}})$. Because its expected payment is at most $\sum_{j=1}^{t}b_{j}$, the user’s expected utility (and thus the social welfare) is positive. This violates global SCP, so the allocation probability $x(v_{0})$ must be positive.
Let $u_{v}$ denote the user’s utility when there is only one user submitting its true value $v$. By Myerson’s lemma, it must be $u_{v}=\int_{0}^{v}x(t)dt$. Because $x(v_{0})>0$, we can find a sufficiently large value $v_{1}$ such that $u_{v_{1}}/v_{1}>x(v_{0})/2$. Moreover, because $x(v)$ is non-decreasing, we have $u_{v}/v\geq u_{v_{1}}/v_{1}$ for any $v\geq v_{1}$.
Next, by weak symmetry, when $n$ users all bid $v$, the social welfare is $\mathsf{SW}(\mathbf{v}_{n})$ regardless of the users’ identities. Thus, by Lemma B.7, for any user identity $i$, we have $\mathsf{SW}(\mathbf{v}_{n})=\mathsf{SW}_{-i}(\mathbf{v}_{n+1})$. Because $\mathsf{SW}(\mathbf{v}_{n+1})=\mathsf{SW}_{-i}(\mathbf{v}_{n+1})+\mathsf{util}_{i}(\mathbf{v}_{n+1})$, we have $\mathsf{util}_{i}(\mathbf{v}_{n+1})=\mathsf{SW}(\mathbf{v}_{n+1})-\mathsf{SW}(\mathbf{v}_{n})$. In other words, given any $v$ and $n$, every user’s utility is the same regardless of the identity. Given any $v$ and $n$, the sum of all users’ utilities is $\mathsf{SW}(\mathbf{v}_{n})-\mu(\mathbf{v}_{n})$, so a user’s utility is $\mathsf{util}(\mathbf{v}_{n})=\frac{1}{n}[\mathsf{SW}(\mathbf{v}_{n})-\mu(\mathbf{v}_{n})]$. Therefore, we have
$\mathsf{SW}(\mathbf{v}_{n+1})=\frac{n+1}{n}\cdot\mathsf{SW}(\mathbf{v}_{n})-\frac{1}{n}\cdot\mu(\mathbf{v}_{n+1}).$ (6)
By Lemma B.5, for any natural number $N$, we can find a sufficiently large value $v\geq v_{1}$ such that $\mu_{n+1}(v)/v<x(v_{0})/4$ for all $n+1\leq N$. Because $u_{v}/v\geq u_{v_{1}}/v_{1}$ for any $v\geq v_{1}$, we have
$\frac{u_{v}}{v}\geq\frac{u_{v_{1}}}{v_{1}}>\frac{x(v_{0})}{2}>\frac{2\mu_{n+1}(v)}{v}.$
Thus, we have $\mu_{n+1}(v)<u_{v}/2$ for all $n+1\leq N$, and Equation 6 implies $\mathsf{SW}(\mathbf{v}_{n+1})\geq\frac{n+1}{n}\cdot\mathsf{SW}(\mathbf{v}_{n})-\frac{1}{2n}\cdot u_{v}$. When there is only one user submitting its true value $v$, the social welfare $\mathsf{SW}(\mathbf{v}_{1})$ is at least $u_{v}$. By the proof of induction, it is easy to see that $\mathsf{SW}(\mathbf{v}_{n})\geq n\cdot u_{v}/2$. Because the block size $k$ is finite, the social welfare $\mathsf{SW}(\mathbf{v}_{n})$ must be upper bounded by $kv$. Thus, we have $u_{v}/v\leq 2k/n$ for all $n\leq N$. However, as we have shown, for a sufficiently large value $v_{1}$, we have $u_{v}/v\geq u_{v_{1}}/v_{1}>x(v_{0})/2$. Consequently, if we choose $N$ such that $N>4k/x(v_{0})$, we have $u_{v}/v\leq 2k/N<x(v_{0})/2$, which leads to a contradiction.